# Influencer Identification on Link Predicted Graphs

### v2

@@ II.4 Heuristic Top-$k$ Algorithms
Transitioning from the single influencer concept to this more practical problem then introduces more complexities. In particular, from Lü et al. listofidentificationtypes , the straightforward approach of taking the highest $k$ nodes by score in a given centrality metric could lead to redundant influences. When information travels relatively unresisted, as in simple contagion (see Table 1), redundancy is often undesirable, as when a single successful interaction suffices to influence an individual, having too much overlap between the chosen nodes’ targets potentially undermines their collective effectiveness and influence. However, in complex contagion, social reinforcement is actually necessary to induce influence cascades, which leads to interesting questions around the straightforward top-$k$ algorithm. We shall thus assess whether the perceived redundancy truly implies reduced performance.
@@
LIR. To reduce the number of adjacent nodes in the chosen set of $k$, Liu et al. lir_original created the local index rank (LIR) algorithm. They first defined the function $Q(x)$ to be $1$ for $x>0$ and 00 otherwise, then set the $LI$ score of each node $u$ as
@@
LIR-2. Inspired by the LIR method described before, Tao et al. lir2 created an LIR-2 method as a quasi-local extension to second-order neighbors. They define the LIR-2 score of a node $u$ to be
@@
Joint Nomination. Dong et al. jointnomination proposed the joint nomination algorithm as an algorithm geared towards immunization purposes, where it seeks to select vertices whose removal can help stop influence spread. Their method spans $k$ rounds. In each round, a random node is selected as a nominator, one of its neighbors is randomly selected as conominator, and finally a random common neighbors of the nominator and conominator is selected into the chosen set, as shown in Algorithm 5. For developing our Social Sphere model we shall amend the method of random selection by including probabilities based on centrality metric score and weights  [^2] In this paper, we deviate from the original version of joint nomination by choosing nodes with probability proportional to centrality score, choosing co-nominators with probability proportional to the weight of the edges between them, and choosing a common neighbor (nominee) $i$ with probability $\frac{w(u,i)\cdot w(v,i)}{\sum_{j\in N(u)\cap N(v)}w(u,j)\cdot w(v,j)}$..
Algorithm 5 Joint Nomination jointnomination
@@
VoteRank. From a different perspective, Zhang et al. voterank proposed VoteRank as an algorithm where influential nodes are chosen by “votes” of the vertices in the graph. There are $k$ rounds of voting. Each node begins with voting power $1$ and score 00 and will divide their voting power among their neighbors, thus changing their scores. In this setting, the node with maximum score is chosen in each round and a resulting penalty given to its neighbors (see Algorithm 6).
Algorithm 6 VoteRank Algorithm voterank
@@
Later on, researchers such as Sun et al. wvoterank have modified VoteRank to fit weighted graphs by changing the score to include the importance of having many neighbors. Researchers such as Li et al. dilvoterank and Kumar and Panda kcorevoterank have combined VoteRank with metrics such as the DIL method (proposed by Liu et al. dil ) and coreness; inspired by these works, we give a generalization of VoteRank for a given centrality metric $c$ (such as the ones in Section II.3) by setting each node’s voting ability to its score with respect to $c$.
Graph Coloring. Finally, aiming to avoid redundancy, Zhao et al. graphcoloring proposed the graph coloring method as a means of selecting separated vertices. They use the Welsh-Powell algorithm to separate the graph into sets such that each set has a unique color and no nodes in the same set share an edge. For a given centrality metric, they then pick the $k$ nodes with highest score from the largest independent set as their chosen $k$ vertices.
Algorithm 7 Welsh-Powell Algorithm graphcoloring
@@
III The Social Sphere Model
@@
III.1 Mathematical Influencers
As mentioned before, one may study social networks through their graph representation, where vertices represent individuals, and their relations are reflected within the edges. In influence maximization, influencers are often defined as those most important to a network, whether structurally or influentially survey2 . In this paper, we shall use a more specific definition of influencer in relation to centrality and influence maximization algorithms.
@@ III.1 Mathematical Influencers
III.2 Future Network Prediction
@@
We shall combine these and look for future influencers by first predicting the state of the future network, like Yanchenko et al. exante did. However, when considering the formation of the future graph, rather than selecting the top edge pairs, we simply take normalized similarity scores to be probabilities of transmission and add them to the graph as weights, using expected value to calculate the distance. Moreover, given the unit of time $t$ we are predicting for, we can find a better predictor for the transmission probability.
@@
III.3 Future Top-$k$ Nodes Prediction
@@
IV Methods
In order to illustrate the different uses of our model and our algorithm, we shall test our algorithm on both a random graph and an empirical dataset: given a graph $G$, we form a training graph from a random subset of existing edges of given size, predict the future state of the training graph, and identify influencers on those predicted graphs. For each graph and parameter tested, we repeated the process ten times, averaging the data. We modelled a dynamic network by taking the training graph as the current configuration and the original graph $G$ as the future network, thus allowing us to assess the efficacy of the influencers found in terms of their ability to “infect” nodes over time when set as the initial spreaders in $G$.
IV.1 Evaluation Methods
@@
To measure the effectiveness mentioned in (I), we quantify influence as the percentage of nodes infected at each unit of time when the chosen influencers are set as initial spreaders on $G$ and an influence model (see Section IV.2) is run. The results are then presented visually in the form of line graphs.
From a different perspective, we describe the predictive success mentioned in (II) as the similarity between our predicted influencers (and their influence) and the influencers found on $G$ (the “true” influencers), quantified in two different ways:
@@ IV.2 Contagion Models
Simple Contagion. In general, simple contagion can be understood as an “infection” of information, where an uninfected individual $j$ becomes infected with probability $p$ once news is passed along to them from an infected individual $i$, in a simple percolation model. For simplicity’s sake, we shall take $p=1$ (following SRI models). Then consider our definition of distance as the expected units of time transmission takes; with this model, influence thus can be approximated as transmitting across each edge $uv$ in $d(u,v)$ units of time. Thus, we can see the following:
@@
(see e.g. Shakarian et al. icmltm for further details). While resistance often varies depending on the individual, in this paper we set it to be a fixed percentage of each vertex’s strength, as considering how perceived plurality can lead to adoption, individuals with larger neighborhoods may have more resistance and similarly the converse may happen.
@@
IV.3 Experimental Setup
We tested our algorithm on an Erdös-Rényi random graph and an existing collaboration network dataset. Erdös and Rényi erdosrenyirgraph defined these random graphs by taking two parameters, $n\in\mathbb{Z}$ and $p\in[0,1]$. The graph is created with $n$ vertices, and each pair of vertices is connected by an edge with probability $p$. We used a graph with $n=500$ and $p=0.05$. The empirical network we used is the arXiv General Relativity and Quantum Cosmology collaboration network from a paper by Leskovec et al. cagrqc , found in the Stanford Large Network Dataset Collection snapnets . The network contains $5242$ nodes and $14496$ edges. In the random graph case, we took $k=5$ and in the arXiv dataset, we took $k=25$. Additionally, we tested the influences of single influencers in each scenario as well.
@@ IV.3 Experimental Setup
V Results
We shall describe here our results obtained though the methods appearing in the previous section, with insightful findings around the efficacy of our model in accurately predicting future influencers and analysis of the comparative performances of various similarity measures, centrality measures, and algorithms. We shall also break down these results and provide a clearer understanding of our model’s capabilities and limitations below.
V.1 Results on a Random Graph
@@
  • the average MSE for each of our algorithms is low, indicating a high similarity to the performance of influencers chosen on the “future” graph,
@@
Algorithm
Common Neighbors
Jaccard
Local Path
Quasi-Local RA
Quasi-Local RA-2
RA-2
Resource Allocation
Overall
Centrality VoteRank
0.32
0.3467
0.3183
0.3567
0.345
0.345
0.3533
0.3407
Graph Coloring
0.0167 0.0167
0.01
0.01
0.01
0.0167 0.0167
0.01381
Joint Nomination
0.0117
0.0183
0.0117
0.0183
0.0133
0.025
0.0117
0.0157
k Highest
0.4567
0.4617
0.4183
0.4267
0.4417
0.4833
0.465
0.4505
LIR
0.2833
0.3267
0.28
0.3367
0.395
0.3633
0.3
0.3264
LIR-2
0.34
0.36
0.34
0.38
0.44 0.44
0.38
0.3829
Single Influencer
0.375
0.3583
0.3583
0.35
0.3583
0.3667
0.375
0.3631
Random
0
0
0.01 0.01 0.01 0.01 0.01
0.0071
VoteRank
0.3231
0.3477
0.3215
0.3569
0.3446
0.3446
0.3538
0.3418
Overall
0.2529
0.2646
0.2439
0.2634
0.2805
0.2855
0.2667
0.2654
@@
Algorithm
Common Neighbors
Jaccard
Local Path
Quasi-Local RA
Quasi-Local RA-2
RA-2
Resource Allocation
Overall
Graph Coloring
0.0233
0.0233
0.0600 0.0600 0.0600
0.0233
0.0233
0.0390
Joint Nomination
0.0067
0.0117
0.0100
0.0133
0.0150
0.0133
0.0100
0.0114
k Highest
0.6867
0.6783
0.5867
0.5833
0.5850
0.6867
0.6850
0.6417
LIR
0.5217
0.5367
0.5067
0.5433
0.6033 0.6033
0.5483
0.5519
LIR-2
0.6400
0.6400
0.6400
0.6600 0.6600 0.6600 0.6600
0.6514
Single Influencer
0.6167
0.6000
0.5583
0.5417
0.5417
0.6000
0.5917
0.5786
Random
0.0000
0.0000
0.0100 0.0100 0.0100 0.0100
0.0000
0.0057
VoteRank
0.6
0.62
0.5954
0.6354
0.6308
0.62
0.6277
0.6185
Overall
0.4338
0.4361
0.4149
0.4264
0.4345
0.4497
0.4411
0.4338
@@
When considering different prediction metrics, we will look only at data from the complex contagion models (see Figure 4 and Figure 5), as the simple contagion data is very close together for each pair of algorithms and prediction metrics, making it less interesting. In most of the graphs, we can see that compared to ‘None,’ the training graph itself, there are a substantial number of predicted graphs that perform consistently better (RA22-2- 2 and QR22-2- 2, for instance). Moreover, the predicted graphs all display influence approaching or, in 4(c), even surpassing the averages for the original graph, suggesting an improvement in prediction when our Social Sphere Model is used.

### v3

@@ 2.4 Heuristic Top-$k$ Algorithms
Transitioning from the single influencer concept to this more practical problem then introduces more complexities. In particular, from Lü et al. [42], the straightforward approach of taking the highest $k$ nodes by score in a given centrality metric could lead to redundant influences. When information travels relatively unresisted, as in simple contagion (see Table 1), redundancy is often undesirable, as when a single successful interaction suffices to influence an individual, having too much overlap between the chosen nodes’ targets potentially undermines their collective effectiveness and influence. However, in complex contagion, social reinforcement is actually necessary to induce influence cascades, which leads to interesting questions around the straightforward top-$k$ algorithm. We shall thus assess whether the perceived redundancy truly implies reduced performance.
@@
LIR. To reduce the number of adjacent nodes in the chosen set of $k$, Liu et al. [38] created the local index rank (LIR) algorithm. They first defined the function $Q(x)$ to be $1$ for $x>0$ and 00 otherwise, then set the $LI$ score of each node $u$ as
@@
LIR-2. Inspired by the LIR method described before, Tao et al. [57] created an LIR-2 method as a quasi-local extension to second-order neighbors. They define the LIR-2 score of a node $u$ to be
@@
Joint Nomination. Dong et al. [17] proposed the joint nomination algorithm as an algorithm geared towards immunization purposes, where it seeks to select vertices whose removal can help stop influence spread. Their method spans $k$ rounds. In each round, a random node is selected as a nominator, one of its neighbors is randomly selected as conominator, and finally a random common neighbors of the nominator and conominator is selected into the chosen set, as shown in Algorithm 5. For developing our Social Sphere model we shall amend the method of random selection by including probabilities based on centrality metric score and weights  [^2] In this paper, we deviate from the original version of joint nomination by choosing nodes with probability proportional to centrality score, choosing co-nominators with probability proportional to the weight of the edges between them, and choosing a common neighbor (nominee) $i$ with probability $\frac{w(u,i)\cdot w(v,i)}{\sum_{j\in N(u)\cap N(v)}w(u,j)\cdot w(v,j)}$..
Algorithm 5 Joint Nomination [17]
@@
VoteRank. From a different perspective, Zhang et al. [66] proposed VoteRank as an algorithm where influential nodes are chosen by “votes" of the vertices in the graph. There are $k$ rounds of voting. Each node begins with voting power $1$ and score 00 and will divide their voting power among their neighbors, thus changing their scores. In this setting, the node with maximum score is chosen in each round and a resulting penalty given to its neighbors (see Algorithm 6).
Algorithm 6 VoteRank Algorithm [66]
@@
Later on, researchers such as Sun et al. [56] have modified VoteRank to fit weighted graphs by changing the score to include the importance of having many neighbors. Researchers such as Li et al. [36] and Kumar and Panda [34] have combined VoteRank with metrics such as the DIL method (proposed by Liu et al. [39]) and coreness; inspired by these works, we give a generalization of VoteRank for a given centrality metric $c$ (such as the ones in Section 2.3) by setting each node’s voting ability to its score with respect to $c$.
Graph Coloring. Finally, aiming to avoid redundancy, Zhao et al. [69] proposed the graph coloring method as a means of selecting separated vertices. They use the Welsh-Powell algorithm to separate the graph into sets such that each set has a unique color and no nodes in the same set share an edge. For a given centrality metric, they then pick the $k$ nodes with highest score from the largest independent set as their chosen $k$ vertices.
Algorithm 7 Welsh-Powell Algorithm [69]
@@
2.5 Comparison to GNN-Based and Temporal Deep Learning Approaches
Recent advances in link prediction and influence maximization have made significant use of geometric deep learning (GDL), particularly Graph Neural Networks (GNNs). These models learn node embeddings that capture complex structural and temporal features of evolving networks. GNN-based models such as GCNs, GraphSAGE, and temporal variants like DyGCN, TGAT, and TGN have shown strong performance in tasks like dynamic link prediction and influence estimation [10, 67, 26].
However, these approaches come at a cost as they typically require extensive training on large datasets and GPU resources; their internal decision processes can be difficult to interpret, limiting transparency; and they often require temporal labels or timestamps that may not be available. In contrast, the Social Sphere Model relies on interpretable, path-based link prediction scores and top-k heuristics to identify future influencers. While it may not match GNNs in raw predictive power on large datasets, its strengths lie in:
  • Reduced computational complexity;
  • Transparency and explainability of results;
  • Applicability to small-scale or under-sampled temporal networks.
This trade-off between interpretability and model expressiveness has been acknowledged in the literature [27], with ongoing research attempting to bridge the gap.
In future work, we aim to extend our framework by integrating learned embeddings into the prediction stage or applying our interpretability heuristics as post hoc analyses to GNN-based results.
3 The Social Sphere Model
@@
3.1 Mathematical Influencers
As mentioned before, one may study social networks through their graph representation, where vertices represent individuals, and their relations are reflected within the edges. In influence maximization, influencers are often defined as those most important to a network, whether structurally or influentially [50]. In this paper, we shall use a more specific definition of influencer in relation to centrality and influence maximization algorithms.
@@ 3.1 Mathematical Influencers
3.2 Future Network Prediction
@@
We shall combine these and look for future influencers by first predicting the state of the future network, like Yanchenko et al. [65] did. However, when considering the formation of the future graph, rather than selecting the top edge pairs, we simply take normalized similarity scores to be probabilities of transmission and add them to the graph as weights, using expected value to calculate the distance. Moreover, given the unit of time $t$ we are predicting for, we can find a better predictor for the transmission probability.
@@
3.3 Future Top-$k$ Nodes Prediction
@@
4 Methods
In order to illustrate the different uses of our model and our algorithm, we shall test our algorithm on a random graph: given a graph $G$, we form a training graph from a random subset of existing edges of given size, predict the future state of the training graph, and identify influencers on those predicted graphs. For each graph and parameter tested, we repeated the process ten times, averaging the data. We modelled a dynamic network by taking the training graph as the current configuration and the original graph $G$ as the future network, thus allowing us to assess the efficacy of the influencers found in terms of their ability to “infect” nodes over time when set as the initial spreaders in $G$.
4.1 Evaluation Methods
@@
To measure the effectiveness mentioned in (I), we quantify influence as the percentage of nodes infected at each unit of time when the chosen influencers are set as initial spreaders on $G$ and an influence model (see Section 4.2) is run. The results are then presented visually in the form of line graphs. From a different perspective, we describe the predictive success mentioned in (II) as the similarity between our predicted influencers (and their influence) and the influencers found on $G$ (the “true" influencers), quantified in two different ways:
@@ 4.2 Contagion Models
Simple Contagion. In general, simple contagion can be understood as an “infection" of information, where an uninfected individual $j$ becomes infected with probability $p$ once news is passed along to them from an infected individual $i$, in a simple percolation model. For simplicity’s sake, we shall take $p=1$ (following SRI models). Then consider our definition of distance as the expected units of time transmission takes; with this model, influence thus can be approximated as transmitting across each edge $uv$ in $d(u,v)$ units of time. Thus, we can see the following:
@@
(see e.g. Shakarian et al. [54] for further details). While resistance often varies depending on the individual, in this paper we set it to be a fixed percentage of each vertex’s strength, as considering how perceived plurality can lead to adoption, individuals with larger neighborhoods may have more resistance and similarly the converse may happen.
@@
4.3 Experimental Setup
We tested our algorithm on an Erdös-Rényi random graph and an existing collaboration network dataset. Erdös and Rényi [19] defined these random graphs by taking two parameters, $n\in\mathbb{Z}$ and $p\in[0,1]$. The graph is created with $n$ vertices, and each pair of vertices is connected by an edge with probability $p$. We used a graph with $n=500$ and $p=0.05$. In the random graph case, we took $k=5$. Additionally, we tested the influences of single influencers in each scenario as well.
@@ 4.3 Experimental Setup
5 Results
@@
  • the average MSE for each of our algorithms is low, indicating a high similarity to the performance of influencers chosen on the “future" graph,
@@
Algorithm Common Neighbors Jaccard Local Path Quasi-Local RA Quasi-Local RA-2 RA-2 Resource Allocation Overall
Centrality VoteRank 0.32 0.3467 0.3183 0.3567 0.345 0.345 0.3533 0.3407
Graph Coloring 0.0167 0.0167 0.01 0.01 0.01 0.0167 0.0167 0.01381
Joint Nomination 0.0117 0.0183 0.0117 0.0183 0.0133 0.025 0.0117 0.0157
k Highest 0.4567 0.4617 0.4183 0.4267 0.4417 0.4833 0.465 0.4505
LIR 0.2833 0.3267 0.28 0.3367 0.395 0.3633 0.3 0.3264
LIR-2 0.34 0.36 0.34 0.38 0.44 0.44 0.38 0.3829
Single Influencer 0.375 0.3583 0.3583 0.35 0.3583 0.3667 0.375 0.3631
Random 0 0 0.01 0.01 0.01 0.01 0.01 0.0071
VoteRank 0.3231 0.3477 0.3215 0.3569 0.3446 0.3446 0.3538 0.3418
Overall 0.2529 0.2646 0.2439 0.2634 0.2805 0.2855 0.2667 0.2654
@@
Algorithm Common Neighbors Jaccard Local Path Quasi-Local RA Quasi-Local RA-2 RA-2 Resource Allocation Overall
Graph Coloring 0.0233 0.0233 0.0600 0.0600 0.0600 0.0233 0.0233 0.0390
Joint Nomination 0.0067 0.0117 0.0100 0.0133 0.0150 0.0133 0.0100 0.0114
k Highest 0.6867 0.6783 0.5867 0.5833 0.5850 0.6867 0.6850 0.6417
LIR 0.5217 0.5367 0.5067 0.5433 0.6033 0.6033 0.5483 0.5519
LIR-2 0.6400 0.6400 0.6400 0.6600 0.6600 0.6600 0.6600 0.6514
Single Influencer 0.6167 0.6000 0.5583 0.5417 0.5417 0.6000 0.5917 0.5786
Random 0.0000 0.0000 0.0100 0.0100 0.0100 0.0100 0.0000 0.0057
VoteRank 0.6 0.62 0.5954 0.6354 0.6308 0.62 0.6277 0.6185
Overall 0.4338 0.4361 0.4149 0.4264 0.4345 0.4497 0.4411 0.4338
@@
When considering different prediction metrics, we will look only at data from the complex contagion models (see Figure 4 and Figure 5), as the simple contagion data is very close together for each pair of algorithms and prediction metrics, making it less interesting. In most of the graphs, we can see that compared to ‘None,’ the training graph itself, there are a substantial number of predicted graphs that perform consistently better (RA22-2- 2 and QR22-2- 2, for instance). Moreover, the predicted graphs all display influence approaching or, in 5(c), even surpassing the averages for the original graph, suggesting an improvement in prediction when our Social Sphere Model is used.